# Consistent Validation for Predictive Methods in Spatial Settings

### v1

@@ B.6.5 Grid Prediction Proofs
Proposition B.16 (Packing and Covering Numbers Wainwright 2019, Lemma 5.5).
@@
The second inequality follows from the first by Prop. B.7 and because $1\in T_{2}$. We therefore focus on proving the case $k=1$.
In light of Thm. 5.1, it suffices to show that for these test location,
@@
The set $\{a/g:1\leq a\leq g\}^{d}$ is a $\frac{1}{g+\epsilon}$-packing for any $\epsilon>0$, and so by the first inequality in Prop. B.16 and Prop. B.3
@@
Applying Wainwright (2019, Lemma 5.7, Equation 5.9) and taking the limit as $\epsilon\to 0^{+}$
@@
Appendix C Additional Experimental Details
@@
In this section, we provide additional details about the data, fitting procedures and validation procedures used in Section 6. Code used in experiments is available anonymously at: https://github.com/DavidRBurt/Consistent-Spatial-Validation. All code is implemented in Python3 (Van Rossum & Drake, 2009). Numpy is also heavily used for data generation and array manipulation (Harris et al., 2020).
C.1 Risk Estimation on Synthetic Data
@@
In this section, we provide additional details and figures for our risk estimation experiment on synthetic data presented in Section 6.1. Section C.1.1 describes the process by which we generate both datasets considered and shows an example of the covariate and response spatial fields for each problem (Figs. 6 and 8). Section C.1.2 describes the procedure used to fit the predictive methods to each dataset. Section C.1.3 describes the implementation of the risk estimation procedures we compare. Section C.1.4 describes the metric reported, and Figs. 7 and 9 show the (signed) relative error of each estimator as we vary the amount of validation data available.
C.1.1 Simulation Data Generating Process Details
@@
This is reminiscent of a Chinese restaurant process (Pitman, 2006, Section 3.1), but the weights are increased more slowly, leading to more clusters being formed and less large clusters typically.
@@
The first $1000$ points generated this way are taken to be the training data, and the remaining $N^{\textup{{val}}}$ points generated this way are the validation data. An example of the training and validation data generated through this process are shown in the top left of Fig. 6.
Generation of test data The test data is $\{(-0.5+a/49,-0.5+b/49):0\leq a,b\leq 49\}$. That is, it is a regular grid on $[-0.5,0.5]^{2}$.
@@
A small diagonal term (1e-12) is added to the diagonal of the covariance matrix to avoid numerical linear algebra errors. Sampling is performed using tensorflow probability (Dillon et al., 2017). We generate two covariates spatial processes via this process $\chi=(\chi^{(1)},\chi^{(2)})$.
@@
$\!\!k(S,S^{\prime})\!=0.5\left(1+\tfrac{\sqrt{3}\|S-S^{\prime}\|_{2}}{0.5}\right)\exp\!\left(-\tfrac{\sqrt{3}\|S-S^{\prime}\|_{2}}{0.5}\right)\!+\!\left(1\!+\sqrt{3}\|\chi(S)-\chi(S^{\prime})\|_{2}\right)\exp\!\left(-\sqrt{3}\|\chi(S)-\chi(S^{\prime})\|_{2}\right).$ (105)
@@
Generation of test site The test site is fixed to be the origin. We generate 2500 response values at the origin so that when we compute the empirical risk we expect it to accurately reflect that actual risk.
@@
A small diagonal term (1e-12) is added to the diagonal of the covariance matrix to avoid numerical linear algebra errors. Sampling is performed using tensorflow probability (Dillon et al., 2017). We generate two covariates via this process $X=(X^{(1)},X^{(2)})$. Generation of response Once the sites and covariates have been generated, the response variable is sampled from a Gaussian process with zero mean. The covariance function of the Gaussian is a sum of two, $2$ dimensional isotropic squared exponential kernels:
@@
C.1.2 Model Fitting
We fit a Gaussian process regression model to using only the first covariate $\chi^{(1)}$ to the training data. The prior is taken to be the same as the data generating process, but with (only) $\chi^{(1)}$ in place of $(\chi^{(1)},\chi^{(2)})$ in the second kernel in Eqn. 105 and Eqn. 107 for the two datasets respectively. The mean of the posterior process is used for predictions, and is calculated using GPFlow (Matthews et al., 2017).
C.1.3 Implementation of Risk Estimation
The holdout is implemented by taking an (unweighted) average of the loss on each validation point. Both nearest neighbor methods are implemented using scikit-learn (Pedregosa et al., 2011) with $kd$-trees and Euclidean distance. For $k^{\star}_{T_{2}}$ nearest neighbors, nearest neighbors is performed for all $k$ that are powers of $2$ less than $N^{\textup{{val}}}$, and the value of $k$ with the smallest bound is used for risk estimation. This is done with $\delta=0.1$ and $\Delta=1$. Table 2 and Table 3 show the values of $k$ chosen for grid and point prediction respectively. For the grid prediction task, $k^{\star}_{T_{2}}$ tends to be small, even as the size of the validation set becomes large. This supported by Cor. B.12, since even $1$-nearest neighbor reliably estimates risk in this setting. For the point prediction task, $k^{\star}_{T_{2}}$ grows with $N^{\textup{{val}}}$.
@@ C.1.3 Implementation of Risk Estimation
1 78 83 79 91 90 80
2 19 16 21 9 10 20
4 3 1 0 0 0 0
Table 3: Value of $k^{\star}_{T_{2}}$ chosen by minimizing the bound in the point prediction task. In this task, there is generally a bias-variance trade-off that must be balances. As a result the value of $k$ that minimizes the upper bound increases as the amount of available validation data increases..
@@
16 1 0 0 0 0 0
32 97 27 0 0 0 0
64 2 73 98 24 0 0
128 0 0 2 76 100 11
256 0 0 0 0 0 89
C.1.4 Metrics Reported and Additional Figures
We use truncated, squared loss,
$\ell(a,b)=\min(1.0,(a-b)^{2}),$ (108)
which is bounded by $1.0$. The empirical risk is calculated as in Eqn. 15. In both cases, we have $2500$ test points that are averaged over; as this is a simple Monte Carlo estimate, of a bounded random variable, we expect this to be concentrated around the mean. Formally, we may apply Hoeffding’s inequality to conclude that with probability at least $0.95$
$|\hat{R}^{\chi}_{Q^{\textup{{test}}}}(h)-R^{\chi}_{Q^{\textup{{test}}}}(h)|\leq 1.0\sqrt{\frac{1}{2\times 2500}\log\frac{2}{0.05}}\approx 0.0272.$ (109)
Figs. 7 and 9 show the relative errors of each estimation, calculated as
$\frac{\hat{R}^{\chi}_{Q^{\textup{{test}}}}(h)-\hat{R}(h)}{\hat{R}^{\chi}_{Q^{\textup{{test}}}}(h)}.$ (110)
@@ C.1.4 Metrics Reported and Additional Figures
C.2 Details for Model Selection Experiment
We present additional details of the experiment discusses in Section 6.2.
Data Generation
The data generation is fully described by Eqn. 16.
Model Fitting
We consider two models. The first is a constant predictor that predicts $0.25$. The second is an affine model (a linear model with an intercept) fit by minimizing the mean absolute error from the line to the training points. This is fit using the Scikit-learn quantile regression with the (default) “HiGHS” solver (Huangfu & Hall, 2015).
Estimation of Risk
The validation estimates used are calculated in the same as the synthetic experiments outlined previously in Section C.1.3. We use $\Delta=1$ in the bound when selecting $k^{\star}_{T_{2}}$, even though the absolute value loss used can be larger than $1$. We don’t expect this to have a significant impact on the results, as the upper bound we minimize is already misspecified in a similar way by not using the actual Lipschitz constant of the function. We again use $\delta=0.1$ when selecting $k^{\star}_{T_{2}}$.
C.3 Details of Air Temperature Prediction Experiment
We now present additional details of the air temperature experiment presented in Section 6.3.
C.3.1 Results for Grid Prediction
Table 4 shows the results for holdout, $1$-nearest neighbor and SNN with the test set each grid point in the map that is located in the continental United States. We see good agreement between all three method. This is expect for $1$-nearest neighbor and SNN.
Holdout 1NN SNN
GWR $\mathbf{0.84\pm 0.03}$ $\mathbf{0.83}$ $\mathbf{0.83}$
Spatial GP $0.96\pm 0.04$ $0.92$ $0.92$
C.3.2 Data Sources

### v2

@@ D.6.5 Grid Prediction Proofs
Proposition D.20 (Packing and Covering Numbers Wainwright 2019, Lemma 5.5).
@@
The second inequality follows from the first by Proposition D.12 and because $1\in T_{2}$. We therefore focus on proving the case $k=1$.
In light of Theorem 5.1, it suffices to show that for these test location,
@@
The set $\{a/g:1\leq a\leq g\}^{d}$ is a $\frac{1}{g+\epsilon}$-packing for any $\epsilon>0$, and so by the first inequality in Proposition D.20 and Proposition D.3
@@
Applying Wainwright [2019, Lemma 5.7, Equation 5.9] and taking the limit as $\epsilon\to 0^{+}$
@@
Appendix E Additional Experimental Details
@@
In this section, we provide additional details about the data, fitting procedures and validation procedures used in Section 6. Code used in experiments is available anonymously at: https://github.com/DavidRBurt/Consistent-Spatial-Validation. Code is almost all implemented in Python3 [Van Rossum and Drake, 2009] (with a small amount of r). Numpy is also heavily used for data generation and array manipulation [Harris et al., 2020].
In Section E.1 we give an overview of our method for estimating ground truth test risk in all experiments. In Section E.3 we describe details of the synthetic experiment described in Section 6.1. In Section E.4 we provide additional details on the air temperature data and tasks in Sections 6.2 and 6.5. In Section E.5 we provide additional details on the UK flat price prediction experiment described in Section 6.3, while in Section E.6 we provide additional details for the wind speed prediction task presented in Section 6.4.
E.1 Monte Carlo Estimation of Ground Truth Test Risk
We would like to compute the exact test risk across the $M^{\textup{{test}}}$ test points:
$R_{Q^{\textup{{test}}}}(h):=(1/M^{\textup{{test}}})\sum_{m=1}^{M^{\textup{{test}}}}\mathbb{E}[\ell(Y^{\textup{{test}}}_{m},h^{\chi}(S^{\textup{{test}}}_{m}))\big{|}S^{\textup{{test}}}_{m},\chi].$ (111)
In all our examples where we report test risk, we have access to some sample $(Y^{\textup{{test}}}_{m})_{m=1}^{M^{\textup{{test}}}}$ that we will use to construct an estimator. Our plan is to instead use the empirical test risk $\hat{R}_{Q^{\textup{{test}}}}(h)$ as ground truth:
$\hat{R}_{Q^{\textup{{test}}}}(h):=(1/M^{\textup{{test}}})\sum_{m=1}^{M^{\textup{{test}}}}\ell(Y^{\textup{{test}}}_{m},h^{\chi}(S^{\textup{{test}}}_{m})).$ (112)
We would like to know how far off the empirical test risk is from the exact test risk. To that end, we observe that
$\hat{R}_{Q^{\textup{{test}}}}(h)-R_{Q^{\textup{{test}}}}(h)$ $=(1/M^{\textup{{test}}})\sum_{m=1}^{M^{\textup{{test}}}}Z_{m},\quad\textrm{ where}$ (113)
$Z_{m}$ $:=\ell(Y^{\textup{{test}}}_{m},h^{\chi}(S^{\textup{{test}}}_{m}))-\mathbb{E}[\ell(Y^{\textup{{test}}}_{m},h^{\chi}(S^{\textup{{test}}}_{m}))|S^{\textup{{test}}}_{m},\chi].$ (114)
By construction, if we assume the expectations exist, each random variable $Z_{m}$ has mean zero. We make two additional assumptions. (1) We assume that the $Z_{m}$ are independent. (2) We assume that (almost surely) $\forall m,Z_{m}\in(a,b)$ for finite $a,b\in\mathbb{R}$. If the loss is bounded by $\Delta$, then such an $a,b$ exist satisfying $b-a\leq\Delta$; following Assumption 2.1, we use this bound moving forward.
Under these assumptions, we can apply Hoeffding’s inequality to conclude that for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$|\hat{R}_{Q^{\textup{{test}}}}(h)-R_{Q^{\textup{{test}}}}(h)|\leq\Delta\sqrt{\frac{1}{2M^{\textup{{test}}}}\log\frac{2}{\delta}}.$ (115)
If we are willing to make the two assumptions above, we next show that we can reach (high probability) conclusions about the (true) relative quality of different estimators on a particular task if they pass a check: namely, we check if, for a small $\delta$ (e.g. $\delta=0.05$), the right-hand side of Equation 115 is smaller than twice the difference between how much closer the “good” estimator is to the estimate of ground truth than the “bad” estimate. To see why this check is sufficient, first observe the following two applications of the triangle inequality:
[…]
$|\text{bad}-\widehat{\text{true}}|$ $\leq|\text{bad}-\text{true}|+|\text{true}-\widehat{\text{true}}|.$ (117)
Using these two inequalities, we can write
$|\text{bad}-\text{true}|-|\text{good}-\text{true}|$ $\geq|\text{bad}-\widehat{\text{true}}|-|\text{good}-\widehat{\text{true}}|-2|\text{true}-\widehat{\text{true}}|.$ (118)
Therefore, to conclude
$|\text{bad}-\text{true}|-|\text{good}-\text{true}|\geq 0,$ (119)
it suffices for
$|\text{bad}-\widehat{\text{true}}|-|\text{good}-\widehat{\text{true}}|\geq 2|\text{true}-\widehat{\text{true}}|.$ (120)
Under the earlier assumptions, we see that Equation 120 is implied (with high probability) by
$|\text{bad}-\widehat{\text{true}}|-|\text{good}-\widehat{\text{true}}|\geq 2\Delta\sqrt{\frac{1}{2M^{\textup{{test}}}}\log\frac{2}{\delta}}.$ (121)
When discussing each experiment, we discuss the plausibility of the assumptions needed to make this argument when justifying our estimated ground truth, as well as specific values for $\Delta$ and $M^{\textup{{test}}}$ and the resulting bound.
E.2 Computational Considerations
E.2.1 Computational Complexity of our Method
We focus on the case $\mathcal{S}=[0,1]^{d}$ with nearest-neighbors implented using a $kd$-tree. Computation of the approximate fill distance is already discussed in Section D.5.4 and is shown to be $O(dN^{\textup{{val}}}\log N^{\textup{{val}}})$. Construction of a $kd$ tree on the validation data is also $O(dN^{\textup{{val}}}\log N^{\textup{{val}}})$. Once constructed, finding the nearest neighbor for each test point requires $O(\log N^{\textup{{val}}})$ computations of a $d$-dimensional Euclidean distance, meaning finding the neighbors is $O(dM^{\textup{{test}}}k\log N^{\textup{{val}}})$. Since $O(\log N^{\textup{{val}}})$ values of $k$ are tried in selecting $k$, this leads to a complexity not more than $O(dM^{\textup{{test}}}N^{\textup{{val}}}(\log N^{\textup{{val}}})^{2}).$ We expect further improvements could be made by storing the nearest neighbor set for intermediate values of $k$, but we do not pursue these.
E.2.2 Computational Setup Used
All experiments were run on a CPU cluster with 36 Intel(R) Xeon(R) W-2295 CPU @ 3.00GHz CPUs and a total of 251 GB of system RAM. In all experiments linear algebra operations operations were allowed to be multithreaded, and so at times all 36 CPUs were used, even if fewer than 36 parallel jobs were run.
E.2.3 Computational Cost of Synthetic Experiment
Data Generation.
Generating the point prediction synthetic data takes around 25 minutes using 10 parallel jobs and has a peak memory usage of around 46GB.
Generating the grid prediction synthetic data takes around 33 minutes using 10 parallel jobs and has a peak memory usage of around 71GB.
Running Experiment.
Running the point prediction task takes around 25 minutes using 10 parallel jobs and has a peak memory usage of around 32GB.
Running the grid prediction task takes around 27 minutes using 10 parallel jobs and has a peak memory usage of around 39GB.
E.2.4 Computational Cost of Bootstrapped Air Temperature Experiment
Data Generation
Running the make file to download air station data takes on the order of 30 seconds and not more than 6GB of RAM after some data has been installed manually as described in the README file in the released code. Fitting the model and computing residuals in order to generate bootstrapped datasets takes around 10.5 minutes and has peak memory usage around 95GB.
Running Experiment
Running the bootstrapped air temperature metro prediction task takes around 66 minutes with 20 parallel jobs and has peak memory usage around 225GB.
Running the bootstrapped air temperature grid prediction task takes around 8.5 hours with 3 parallel jobs and has peak memory usage around 150GB.
E.2.5 Computational Cost of UK House Price Experiment
Data Processing
Downloading and processing the data takes around 7 seconds and has a peak memory usage of under 3GB.
Running Experiment
Running the UK House price prediction task takes around 4.6 hours with 5 parallel jobs and has peak memory usage around 70GB.
E.2.6 Computational Cost of Wind Speed
Data Processing
Downloading and processing the data takes around 3.5 minutes and has a peak memory usage of under 12GB.
Running Experiment
Running the wind spped prediction task takes around 8.5 hours with 15 parallel jobs and has peak memory usage around 15GB.
E.2.7 Computational Cost of Real Data Air Temperature Experiment
Time to process the data has been previously described in Section E.2.4.
Running Experiment
Running all the air temperature tasks takes around 10 minutes and has peak memory usage around 9GB.
E.2.8 Computational Cost of Model Selection Experiment
Generating data for and running the synthetic model selection experiment takes under a minute of time and under 4GB of RAM.
E.3 Risk Estimation on Synthetic Data
@@
In this section, we provide additional details and figures for our risk estimation experiment on synthetic data presented in Section 6.1. Section E.3.1 describes the process by which we generate both datasets considered and shows an example of the covariate and response spatial fields for each problem (Figures 4 and 6). Section E.3.3 describes the procedure used to fit the predictive methods to each dataset. Section E.3.4 describes the implementation of the risk estimation procedures we compare. Section E.3.6 describes the metric reported, and Figures 5 and 7 show the (signed) relative error of each estimator as we vary the amount of validation data available.
E.3.1 Simulation Data Generating Process Details
@@
This is reminiscent of a Chinese restaurant process [Pitman, 2006, Section 3.1], but the weights are increased more slowly, leading to more clusters being formed and less large clusters typically.
@@
The first $1000$ points generated this way are taken to be the training data, and the remaining $N^{\textup{{val}}}$ points generated this way are the validation data. An example of the training and validation data generated through this process are shown in the top left of Figure 4.
Generation of test data The test data is $\{(-0.5+a/29,-0.5+b/29):0\leq a,b\leq 49\}$. That is, it is a regular grid on $[-0.5,0.5]^{2}$. We generate 50 values of each response variable on each. grid point.
@@
A small diagonal term (1e-12) is added to the diagonal of the covariance matrix to avoid numerical linear algebra errors. Sampling is performed using Tensorflow probability [Dillon et al., 2017]. We generate two covariates spatial processes via this process $\chi=(\chi^{(1)},\chi^{(2)})$.
@@
$k(S,S^{\prime})$ $=0.5\left(1+\frac{\sqrt{3}\|S-S^{\prime}\|_{2}}{0.5}\right)\exp\!\left(-\frac{\sqrt{3}\|S-S^{\prime}\|_{2}}{0.5}\right)$ (124)
$+\left(1+\sqrt{3}\|\chi(S)-\chi(S^{\prime})\|_{2}\right)\exp\left(-\sqrt{3}\|\chi(S)-\chi(S^{\prime})\|_{2}\right).$ (125)
@@
Generation of test site The test site is fixed to be the origin. We generate 45000 response values at the origin so that when we compute the empirical risk we expect it to accurately reflect that actual risk.
@@
A small diagonal term (1e-12) is added to the diagonal of the covariance matrix to avoid numerical linear algebra errors. Sampling is performed using tensorflow probability [Dillon et al., 2017]. We generate two covariates via this process $X=(X^{(1)},X^{(2)})$.
Generation of response Once the sites and covariates have been generated, the response variable is sampled from a Gaussian process with zero mean. The covariance function of the Gaussian is a sum of two, $2$ dimensional isotropic squared exponential kernels:
@@
E.3.2 Loss Function
We use truncated, squared loss,
$\ell(a,b)=\min(1.0,(a-b)^{2}),$ (128)
which is bounded by $1.0$. The empirical risk is calculated as in Equation 112.
E.3.3 Model Fitting
We fit a Gaussian process regression model to using only the first covariate $\chi^{(1)}$ to the training data. The prior is taken to be the same as the data generating process, but with (only) $\chi^{(1)}$ in place of $(\chi^{(1)},\chi^{(2)})$ in the second kernel in Equation 125 and Equation 127 for the two datasets respectively. The mean of the posterior process is used for predictions, and is calculated using GPFlow [Matthews et al., 2017].
E.3.4 Implementation of Risk Estimation
The holdout is implemented by taking an (unweighted) average of the loss on each validation point. Both nearest neighbor methods are implemented using scikit-learn [Pedregosa et al., 2011] with $kd$-trees and Euclidean distance. For $k^{\star}_{T_{2}}$ nearest neighbors, nearest neighbors is performed for all $k$ that are powers of $2$ less than $N^{\textup{{val}}}$, and the value of $k$ with the smallest bound is used for risk estimation. This is done with $\delta=r$, with $r$ calculated as in Section D.5.4, and $\Delta=1$. Table 2 and Table 3 show the values of $k$ chosen for grid and point prediction respectively. For the grid prediction task, $k^{\star}_{T_{2}}$ tends to be small, even as the size of the validation set becomes large. This supported by Corollary D.16, since even $1$-nearest neighbor reliably estimates risk in this setting. For the point prediction task, $k^{\star}_{T_{2}}$ grows with $N^{\textup{{val}}}$.
@@ E.3.4 Implementation of Risk Estimation
1 82 85 81 82 66 50
2 15 15 19 18 34 41
4 3 0 0 0 0 9
Table 3: Value of $k^{\star}_{T_{2}}$ chosen by minimizing the bound in the point prediction task. In this task, there is generally a bias-variance trade-off that must be balanced. As a result the value of $k$ that minimizes the upper bound increases as the amount of available validation data increases.
@@
16 7 0 0 0 0 0
32 93 33 0 0 0 0
64 0 67 99 14 0 0
128 0 0 1 86 98 0
256 0 0 0 0 2 100
E.3.5 Monte Carlo Estimation of Test Risk
Following the argument in Section E.1, we use the empirical test risk Equation 112 in place of the test risk as ground truth in synthetic experiments. The assumption that $(Z_{m})_{m=1}^{M^{\textup{{test}}}}$ are independent holds by the description of the data generating process, because the $(Y_{m})_{m=1}^{M^{\textup{{test}}}}$ are conditionally independent and $Z_{m}$ is a function of $Y_{m}$. The assumption that $Z_{m}$ is almost surely bounded holds with $\Delta=1$ by our choice of truncated squared loss. Further, in both synthetic experiments, we take $M^{\textup{{test}}}=45000$. Combining these gives that with probability at least $0.95$
$|\hat{R}_{Q^{\textup{{test}}}}(h)-R_{Q^{\textup{{test}}}}(h)|\leq\sqrt{\frac{1}{2\times 45000}\log\frac{2}{0.05}}\leq 0.0065.$ (129)
Figure 1 shows the absolute difference between each estimator and the empirical test risk across 100 seeds. We see that the difference between the estimators is generally larger than twice Equation 129, and so by the argument in Section E.1, we expect our estimate of ground truth to be accurate enough that the difference in performance of the methods is not simply due to error in estimating the ground truth.
E.3.6 Metrics Reported and Additional Figures
Figures 5 and 7 show the relative errors of each estimation, calculated as
$\frac{\hat{R}_{Q^{\textup{{test}}}}(h)-\hat{R}(h)}{\hat{R}_{Q^{\textup{{test}}}}(h)}.$ (130)
@@ E.3.6 Metrics Reported and Additional Figures
E.4 Air Temperature Tasks
We now provide additional details about data source, pre-processing, model fitting and risk estimation for the air temperature dataset. These are identical between the real response experiment and the partially synthetic experiment, except for the bootstrapping procedure described in Section E.4.6. We also provide additional experimental results on a grid prediction task for both the bootstrapped and original datasets.
E.4.1 Data Sources